# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v2

@@ Appendix D Lower Bound of the Learner’s Wealth
Assume an online learner receives a loss function $\ell_{t}(\theta):=\log(1-g_{t}\theta)$ after committing a point $\theta_{t}\in\mathcal{K}_{t}$ in its decision space $\mathcal{K}_{t}$ at $t$. Denote $d_{*}:=\max\limits_{t\geq 1}\lvert d_{t}\rvert$ with $d_{t}\geq|g_{t}|$. Then, if the online learner plays Online Newton Step (Algorithm 2), its wealth satisfies
$W_{T}\gtrsim\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{(\sum_{t=1}^{T}g_{t})^{2}}{\sum_{t=1}^{T}g_{t}^{2}+\lvert\sum_{t=1}^{T}g_{t}\rvert}\right)\bigg{/}\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{\frac{1}{2\gamma}},$ (25)
where the step size $\gamma$ satisfies $\gamma\leq\frac{1}{2}\min\{\frac{d_{t}}{G_{t}},1\}$ with $G_{t}:=\max_{\theta\in\mathcal{K}_{t}}|\nabla\ell_{t}(\theta)|$ denoting the upper bound of the gradient $\nabla\ell_{t}(\theta)$.
@@
$W_{1}$ $=W_{0}(1-g_{1}\theta_{1}).$ (27)
We start with $W_{0}=1$, then by recursion
$W_{T}=W_{0}\cdot\prod_{t=1}^{T}(1-g_{t}\theta_{t})=\prod_{t=1}^{T}(1-g_{t}\theta_{t}),$ (28)
thus we can express $\log(W_{T})$ as:
$\log(W_{T})=\sum_{t=1}^{T}\log(1-g_{t}\theta_{t}).$ (29)
Similarly, when we choose a signed constant $u$ in hindsight,
$\log(W_{T}(u))=\sum_{t=1}^{T}\log(1-g_{t}u)$ (30)
@@
$\log(W_{T}(u))-\log(W_{T})$ $=\sum_{t=1}^{T}\log(1-g_{t}u)-\sum_{t=1}^{T}\log(1-g_{t}\theta_{t})$
$=-\sum_{t=1}^{T}\log(1-g_{t}\theta_{t})-(-\sum_{t=1}^{T}\log(1-g_{t}u))$
$=\sum_{t=1}^{T}-\log(1-g_{t}\theta_{t})-\sum_{t=1}^{T}-\log(1-g_{t}u).$
The equation can be can be interpreted as the regret of an algorithm, where $\theta_{t}$ is played against losses defined by $\ell_{t}(\theta)=-\log(1-g_{t}\theta)$. Suppose $\text{Regret}_{T}(u)$ is the regret of our method, we have
$\log(W_{T})=\log(W_{T}(u))-\text{Regret}_{T}(u),$ (31)
Given that the loss function $\ell_{t}(\theta)=-\log(1-g_{t}\theta)$ is exp-concave by definition, the task of choosing $v_{t}$ is actually an online exp-concave optimization problem. In the previous section, we obtained $\text{Regret}_{T}(u)=O\left(\log\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{\frac{1}{2\gamma}}\right)$ for Algorithm 2. Now, we can use equation (31) to obtain the lower bound for $W_{T}$. Noting that the term $\gamma D_{1}^{2}/2$ in the regret bound (24) is potentially dominated by the first term $\frac{1}{2\gamma}\cdot\log\left(1+\sum_{t=1}^{T}\frac{g_{t}^{2}}{(1-g_{t}\theta_{t})^{2}}\right)$, as the first term grows with $T$, and taking the exponential on both sides of (31) lead to:
$W_{T}\gtrsim\frac{W_{T}(u)}{(\sum_{t=1}^{T}g_{t}^{2})^{\frac{1}{2\gamma}}}\text{ for all }|u|\leq\frac{1}{2d_{*}}.$ (32)
Next, we will demonstrate that a suitable value of $u$ can be found such that the ratio $W_{T}(u)/(\sum_{t=1}^{T}g_{t}^{2})^{\frac{1}{2\gamma}}$ is sufficiently high to assure low regret of Algorithm 2. Consider
$u=\frac{-\sum_{t=1}^{T}g_{t}}{2d_{*}\cdot\left(\sum_{t=1}^{T}g_{t}^{2}+\left\lvert\sum_{t=1}^{T}g_{t}\right\rvert\right)}\in[-\frac{1}{2d_{*}},\frac{1}{2d_{*}}],$
where $d_{*}:=\max\limits_{t}\lvert d_{t}\rvert$, meaning that $d_{*}\geq d_{t}$ for all $t\geq 1$. Since $g_{t}\in[-d_{t},d_{t}]$, $u\in[-1/2d_{*},1/2d_{*}]$, then we have $-g_{t}u\in[-d_{t}/2d_{*},d_{t}/2d_{*}]\subseteq[-1/2,1/2]$.
Define
$g_{t}:=\phi(x_{t})-\phi(y_{t}),\quad S_{T}:=\sum_{t=1}^{T}g_{t},\quad Q_{T}:=\sum_{t=1}^{T}g_{t}^{2}.$ (33)
then based on equation (30) and the tangent bound $\log(1+a)\geq a-a^{2}$ for $a\in[-1/2,1/2]$:
$\log(W_{T}(u))$ $=\sum_{t=1}^{T}\log(1-g_{t}u)$
$\geq-\sum_{t=1}^{T}g_{t}u-\sum_{t=1}^{T}(-g_{t}u)^{2}$
$=-\sum_{t=1}^{T}g_{t}\cdot u-\sum_{t=1}^{T}g_{t}^{2}\cdot(u)^{2}$
$=-S_{T}\cdot\frac{-S_{T}}{2d_{*}Q_{T}+2d_{*}\lvert S_{T}\rvert}-Q_{T}\cdot\left(\frac{-S_{T}}{2d_{*}Q_{T}+2d_{*}\lvert S_{T}\rvert}\right)^{2}$
$=\frac{S_{T}^{2}}{2d_{*}Q_{T}+2d_{*}\lvert S_{T}\rvert}-\frac{Q_{T}}{2d_{*}Q_{T}+2d_{*}\lvert S_{T}\rvert}\cdot\frac{S_{T}^{2}}{2d_{*}Q_{T}+2d_{*}\lvert S_{T}\rvert}$
$=\frac{1}{2d_{*}}\cdot\frac{S_{t}^{2}}{Q_{T}+\lvert S_{T}\rvert}-\frac{1}{4d_{*}^{2}}\cdot\frac{Q_{T}}{Q_{T}+\lvert S_{T}\rvert}\cdot\frac{S_{T}^{2}}{Q_{T}+\lvert S_{T}\rvert}$
$\geq\frac{1}{2d_{*}}\cdot\frac{S_{T}^{2}}{Q_{T}+\lvert S_{T}\rvert}-\frac{1}{4d_{*}^{2}}\cdot\frac{S_{T}^{2}}{Q_{T}+\lvert S_{T}\rvert}\quad(\text{since }\frac{Q_{T}}{Q_{T}+\lvert S_{T}\rvert}\leq 1)$
$=\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{T}^{2}}{Q_{T}+\lvert S_{T}\rvert}.$
According to (32), we get the following bound of wealth at time T:
$W_{T}\gtrsim\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{(\sum_{t=1}^{T}g_{t})^{2}}{\sum_{t=1}^{T}g_{t}^{2}+\lvert\sum_{t=1}^{T}g_{t}\rvert}\right)\bigg{/}\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{\frac{1}{2\gamma}}.$ (34)
@@ Appendix E Proof of Proposition 1
The proof of Proposition 1 and 2 is based on a modification of Chugg et al. [2023]. The key difference is that we have made the exponent $1/2\gamma$ in the regret bound explicit, which plays a crucial role in deriving the expected rejection time of our algorithm Additionally, we extend the range of $g_{t}$ from $[-1,1]$ to an adaptive interval$[-d_{t},d_{t}]$ for each $t$, and provide a more explicit proof of the statistical guarantees for our algorithm. This range of $\theta_{t}$ is necessary for text detection because scores of texts are unknown and and do not have an explicit predefined bound, as mentioned in the first paragraph of Appendix C. We can divide the proof of Proposition 1 into 3 parts as below.
1. Level-$\alpha$ Sequential Test. In Algorithm 1, we treat $\{W_{t}\geq 1/\alpha\text{ or }W_{T}>Z/\alpha\}$ as reject "$H_{0}$". It is a level-$\alpha$ sequential test means that, when $H_{0}$ holds:
$\sup_{P\in H_{0}}P(\exists t\geq 1:W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha)\leq\alpha,\quad\text{or equivalently}\quad\sup_{P\in H_{0}}P(\tau<\infty)\leq\alpha.$ (35)
Previously, we have defined the minimum rejection time as $\tau=\arg\inf_{t}\{W_{t}\geq 1/\alpha\text{ or }W_{T}\geq Z/\alpha\}$, where $Z\sim\text{Unif}(0,1)$.
Proof.
@@
$\mathbb{E}_{P}[W_{t}|\mathcal{F}_{t-1}]=\mathbb{E}_{P}\left[(1-g_{t}\theta_{t})\times W_{t-1}\Bigg{|}\mathcal{F}_{t-1}\right]=W_{t-1}(1-\theta_{t}\cdot\mathbb{E}_{P}[\phi({x}_{t})-\phi({y}_{t})])=W_{t-1},$ (37)
@@
thus $(W_{t})_{t\geq 1}$ is a $P$-martingale with $W_{0}=1$. Since $g_{i}\in[-d_{i},d_{i}]$ and $\theta_{i}\in[-1/2d_{i},1/2d_{i}]$, we have $g_{i}\theta_{i}\in[-1/2,1/2]$ for all $t$, then $W_{t}=\prod_{i=1}^{t}(1-g_{i}\theta_{i})$ remains non-negative for all $t$. Thus, we can apply Ville’s inequality [Ville, 1939] to estabilish that $P(\exists t\geq 1:W_{t}\geq 1/\alpha)\leq\alpha$. This inequality shows that the sequential test: “reject $H_{0}$ once the wealth $W_{t}$ reaches $1/\alpha$" maintains a level-$\alpha$ type-I error rate. If there exists a time budget $T$, we will verify the final step $W_{T}\geq Z/\alpha$ of the algorithm, which is validated by the randomized Ville’s inequality of Ramdas and Manole [2023]. ∎
@@
2. Asymptotic power one. Test $\phi$ has asymptotic power $\beta=1$ means that when $H_{1}$ ($\mu_{x}\neq\mu_{y}$) holds, our algorithm will ensure that wealth $W_{t}\geq 1/\alpha$ in finite time $t$ to reject $H_{0}$, that is:
$\sup_{P\in H_{1}}P(\tau=\infty)\leq 1-\beta=0.$ (38)
In Appendix D, we get the following guarantee on $W_{t}$, with $W_{0}=1$:
$W_{T}\gtrsim\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{\left(\sum_{t=1}^{T}g_{t}\right)^{2}}{\sum_{t=1}^{T}g_{t}^{2}+\left\lvert\sum_{t=1}^{T}g_{t}\right\rvert}\right)\bigg{/}\left(\sum_{t=1}^{T}g_{t}^{2}\right)^{\frac{1}{2\gamma}}.$ (39)
According to our definitions: $S_{t}=\sum_{i=1}^{t}g_{i}$, $Q_{t}=\sum_{i=1}^{t}g_{i}^{2}$, and $\lvert g_{i}\rvert\leq d_{i}$, where $d_{*}\geq d_{i}$. By the inequality (39), we can derive:
$W_{t}\gtrsim\frac{1}{Q_{t}^{\frac{1}{2\gamma}}}\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{Q_{t}+|S_{t}|}\right)\geq\frac{1}{(td_{*}^{2})^{\frac{1}{2\gamma}}}\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right),\quad\forall t\geq 1$ (40)
By definition of the rejection time, we know $\{\tau>t\}\subseteq\{W_{t}<\frac{1}{\alpha}\}$. Because $\{\tau=\infty\}\subseteq\{\tau\geq t\}$ for all $t\geq 1$, we can ultimately get $P(\tau=\infty)\leq\liminf_{t\to\infty}P(\tau>t)\leq\liminf_{t\to\infty}P(W_{t}<1/\alpha)$. By the second inequality of (40),
$P(W_{t}<1/\alpha)$ $\lesssim P\left(\frac{1}{(td_{*}^{2})^{\frac{1}{2\gamma}}}\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right)<1/\alpha\right)$
$=P\left(\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right)<(td_{*}^{2})^{\frac{1}{2\gamma}}/\alpha\right)$
$\leq P\left(-\sqrt{\frac{4d_{*}^{4}+4d_{*}^{3}}{2d_{*}-1}\cdot\frac{\log(t^{\frac{1}{2\gamma}}d_{*}^{\frac{1}{\gamma}}\alpha)}{t}}<\frac{S_{t}}{t}<\sqrt{\frac{4d_{*}^{4}+4d_{*}^{3}}{2d_{*}-1}\cdot\frac{\log(t^{\frac{1}{2\gamma}}d_{*}^{\frac{1}{\gamma}}/\alpha)}{t}}\right).$
It is almost surely that $S_{t}/t=\frac{1}{t}\sum_{i=1}^{t}(\phi({x}_{t})-\phi({y}_{t}))$ converges to $(\mu_{x}-\mu_{y})$ as $t\rightarrow\infty$, according to the Strong Law of Large Numbers. We recall that under $H_{1}:\mu_{x}-\mu_{y}\neq 0$. On the other hand, $\frac{4d_{*}^{4}+4d_{*}^{3}}{2d_{*}-1}\cdot\frac{\log(t^{\frac{1}{2\gamma}}d_{*}^{\frac{1}{\gamma}}/\alpha)}{t}\to 0$ as $t\rightarrow\infty$. Thus, if we let $a_{t}$ be the event that $\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right)<(td_{*}^{2})^{\frac{1}{2\gamma}}/\alpha$ , we see that $\mathcal{1}_{a_{t}}\to 0$ almost surely. By the dominated convergence theorem,
[…]
$d_{*}$ is the largest absolute difference between two scores $\phi({x}_{t})$ and $\phi({y}_{t})$ for all $t$, thus, $d_{*}>0.5$ can always be guaranteed. This completes the argument of asymptotic power one.
3. Expected stopping time. When there is no constraint on time budget $T$ and under the assumption that $H_{1}$ is true, we have

### v3

@@ Appendix D Lower Bound on the Learner’s Log-Wealth
Assume an online learner receives a loss function $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$ after committing a point $\theta_{t}\in\mathcal{K}_{t}$ in its decision space $\mathcal{K}_{t}$ at time $t$. Denote $d_{*}:=\max_{t\geq 1}d_{t}$ with $d_{t}\geq|g_{t}|$. Then, if the online learner plays Online Newton Step (Algorithm 2), for any benchmark $u\in K_{*}:=\bigcap_{t=1}^{T}\,\mathcal{K}_{t}$, the value of the wealth satisfies
$\ln W_{T}$ $=\ln W_{T}(u)-\mathrm{Regret}_{T}(\mathrm{ONS})$
$\geq\ln W_{T}(u)-\left(\frac{1}{2\gamma}\cdot\ln\left(1+4d_{*}^{2}T\right)+\frac{\gamma}{2}(\theta_{1}-u)^{2}\right),$
where we define the wealth at time $T$ as $W_{T}(\theta):=\prod_{t=1}^{T}(1-g_{t}\theta)$, $\theta_{1}\in\mathcal{K}_{*}$ is the initial point of ONS, and the inverse of the step size $\gamma$ satisfies $\gamma=\frac{1}{2}\min\{\frac{d_{t}}{G_{t}},1\}$, with $G_{t}:=\max_{\theta\in\mathcal{K}_{t}}|\nabla\ell_{t}(\theta)|$ denoting the upper bound of the gradient $\nabla\ell_{t}(\theta)$.
@@
$W_{T}$ $=W_{T-1}(1-g_{T}\theta_{T}).$
We start with $W_{0}=1$, then by recursion,
$W_{T}=W_{0}\cdot\prod_{t=1}^{T}(1-g_{t}\theta_{t})=\prod_{t=1}^{T}(1-g_{t}\theta_{t}),$
thus we can express $\ln(W_{T})$ as:
$\ln(W_{T})=\sum_{t=1}^{T}\ln(1-g_{t}\theta_{t}).$ (16)
Similarly, the benchmark $u$ has:
$\ln(W_{T}(u))=\sum_{t=1}^{T}\ln(1-g_{t}u).$ (17)
@@
$\ln(W_{T}(u))-\ln(W_{T})$ $=\sum_{t=1}^{T}\ln(1-g_{t}u)-\sum_{t=1}^{T}\ln(1-g_{t}\theta_{t})$
$=-\sum_{t=1}^{T}\ln(1-g_{t}\theta_{t})-(-\sum_{t=1}^{T}\ln(1-g_{t}u))$
$=\sum_{t=1}^{T}-\ln(1-g_{t}\theta_{t})-\sum_{t=1}^{T}-\ln(1-g_{t}u).$
The above can be connected to the regret of the online learner, where the loss function is defined as $\ell_{t}(\theta):=-\ln(1-g_{t}\theta)$. Consequently, we have
$\ln(W_{T})=\ln(W_{T}(u))-\text{Regret}_{T}(u).$ (18)
Now applying the regret bound equation 15 leads to the result. ∎
@@ Appendix E Proof of Proposition 1
Some of the ideas in the proof is adapted from Dai et al. [2025], with the key difference that we consider a two-sided test, whereas their analysis focuses on a one-sided test with Bernoulli random variables. Additionally, we consider the range of $g_{t}$ within a time-varying interval $[-d_{t},d_{t}]$ for each round $t$. We can divide the proof of Proposition 1 into three parts as below.
1. Level-$\alpha$ Sequential Test. In Algorithm 1, We treat the event $\{W_{t}\geq 1/\alpha\}$ as rejecting the null hypothesis $H_{0}$. A level-$\alpha$ sequential test is one that guarantees, under $H_{0}$, the probability of false rejection is at most $\alpha$, that is.
$\sup_{P\in H_{0}}P(\exists t\geq 1:W_{t}\geq 1/\alpha)\leq\alpha,\quad\text{orequivalently}\quad\sup_{P\in H_{0}}P(\tau<\infty)\leq\alpha.$
@@
$\mathbb{E}_{P}[W_{t}|\mathcal{F}_{t-1}]=\mathbb{E}_{P}\left[(1-g_{t}\theta_{t})\times W_{t-1}\Bigg{|}\mathcal{F}_{t-1}\right]=\mathbb{E}_{P}\left[W_{t-1}\cdot(1-\theta_{t}\cdot(\phi({x}_{t})-\phi({y}_{t})))\Bigg{|}\mathcal{F}_{t-1}\right]=W_{t-1},$
where the last equality is because $\theta_{t}$ is fully determined by $\mathcal{F}_{t-1}$, and hence, given $\mathcal{F}_{t-1}$, $\theta_{t}$ is independent of $\left(\phi({x}_{t})-\phi({y}_{t})\right)$, where the latter has conditional expectation zero under $H_{0}$.
@@
Thus, $(W_{t})_{t\geq 1}$ is a $P$-martingale with $W_{0}=1$. Given that $g_{i}\in[-d_{i},d_{i}]$ and $\theta_{i}\in[-1/2d_{i},1/2d_{i}]$, we have $g_{i}\theta_{i}\in[-1/2,1/2]$ for all $t$, then $W_{t}=\prod_{i=1}^{t}(1-g_{i}\theta_{i})$ remains non-negative for all $t$. Therefore, we can apply Ville’s inequality [Ville, 1939] to establish that $P(\exists t\geq 1:W_{t}\geq 1/\alpha)\leq\alpha$. This inequality shows that the sequential test: “reject $H_{0}$ once the wealth $W_{t}$ reaches $1/\alpha$" controls the type-I error at level $\alpha$. If there exists a time budget $T$, we will verify the final step $W_{T}\geq Z/\alpha$ of the algorithm, i.e., we treat $\{W_{t}\geq 1/\alpha\text{ or }W_{T}>Z/\alpha\}$ as reject "$H_{0}$", which is validated by the randomized Ville’s inequality of Ramdas and Manole [2023].
@@
2. Asymptotic Power One. Test $\phi$ has asymptotic power $\beta=1$ means that when $H_{1}$ (i.e., $\mu_{x}\neq\mu_{y}$) holds, our algorithm will ensure that wealth $W_{t}\geq 1/\alpha$ in finite time $t$ to reject $H_{0}$, that is:
$\sup_{P\in H_{1}}P(\tau=\infty)\leq 1-\beta=0.$
We first recall the relation between regret and wealth in the testing-by-betting game (c.f. equation 18), i.e.,
$\ln(W_{t})=\ln(W_{t}(\theta_{*}))-\text{Regret}_{t}(\theta_{*}),$ (20)
where $W_{t}$ is the learner’s wealth at round $t$, $\theta_{*}$ is any benchmark which lies in the decision space $\mathcal{K}_{*}=\left[-\frac{1}{2d_{*}},\frac{1}{2d_{*}}\right]$.
Denote $\omega_{*}:=\mathbb{E}[\ln(1-g\theta_{*})]$ the expected wealth of the benchmark in a single round. Taking the expectation on both sides of equation 20, we have
$\mathbb{E}[\ln(W_{t})]$ $=\mathbb{E}[\ln(W_{t}(\theta_{*}))]-\mathbb{E}\left[\text{Regret}_{t}(\theta_{*})\right]$
$=\mathbb{E}\left[\sum_{s=1}^{t}\ln(1-g_{s}\theta_{*})\right]-\mathbb{E}[\text{Regret}_{t}(\theta_{*})]$
$=t\omega_{*}-\mathbb{E}\left[\text{Regret}_{t}(\theta_{*})\right],$
where the last equality holds by assuming that the random variables $(g_{s})_{s\geq 1}$ are i.i.d.
We now analyze the probability that the null has not been rejected by time $t$, which is when the event $\{W_{t}<\frac{1}{\alpha}\}$ holds. We have
$\mathbb{P}\left[W_{t}<\frac{1}{\alpha}\right]$ $=\mathbb{P}\left[\ln(W_{t})<\ln\left(\frac{1}{\alpha}\right)\right]$
$=\mathbb{P}\left[\ln(W_{t})-\mathbb{E}[\ln(W_{t}(\theta_{*}))]<\ln\left(\frac{1}{\alpha}\right)-\mathbb{E}[\ln(W_{t}(\theta_{*}))]\right]$